A number increased by 20% is 16. Find the number.
Found 2 solutions by checkley71, BGluvsmath:
Answer by checkley71(8403)

X+(.2X)=16
1.2X=16
X=16/1.2
X=4/.3
X=13 1/3
PROOF
13 1/3+.2(1 1/3)=16
40/3+.2*4/3=16
40/3+8/3=16
48/3=16
16=16

Answer by BGluvsmath(64)

x+.2x=16
1.2x=16
x=13 1/3